How do you write a function that has x-intercepts of -3 and 4 when graphed?
Answer by Alan3354(69443)

How do you write a function that has x-intercepts of -3 and 4 when graphed?
------------
The simplest is f(x) = (x+3)*(x-4)
Any number of multiplying factors can be added.
eg, f(x) = x*(x+3)*(x-4), f(x) = (x+k)*(x+3)*(x-4) where k is any constant.